OpenStudy (anonymous):

Mike and Kim invest $12,000 in equipment to print yearbooks for schools. Each yearbook costs $5 to print and sells for $15. How many yearbooks must they sell before their business breaks even? A. 800 yearbooks B. 1,200 yearbooks C. 2400 yearbooks D. 1,800 yearbooks

OpenStudy (anonymous):

c

OpenStudy (anonymous):

@skilyrboren I think it's D prove you're right.

OpenStudy (anonymous):

12,000 devided by 5=2400. My tutor taught me that..maybe she is wrong!? D;

OpenStudy (anonymous):

The Selling price of each book is 15 so to come on break even point , multiply 15 X 800 =1200.. so they will reach on break even that is : no profit nor loss

OpenStudy (anonymous):

sorry that was 12000 not 1200 .

OpenStudy (anonymous):

I actually got B I just picked D so I can make you show your work.

OpenStudy (anonymous):

thanks asad that was the best answer

OpenStudy (anonymous):

no problem :)

OpenStudy (kropot72):

Let n = number of yearbooks to be printed. Expenses = 12000 + 5n Income = 15n To break even the income must equal the expenses. 12000 + 5n =15n 10n = 12000 n = 1200